A train left Dhaka and traveled towards Chittagong. Five hours later another train also left for Chittagong, traveling at 40 miles per hour with a goal of catching up with the first train bound for Chittagong. The second train traveled for three hours and caught up to the first. How fast was the first train going?

Answer by stanbon(75887)

A train left Dhaka and traveled towards Chittagong. Five hours later another train also left for Chittagong, traveling at 40 miles per hour with a goal of catching up with the first train bound for Chittagong. The second train traveled for three hours and caught up to the first. How fast was the first train going?
---------------------
1st train DATA:
time = 8 hrs ; rate = r mph ; distance = r*t = 8r miles
-------
2nd train DATA:
time = 3 hrs ; rate = 40 mph ; distance = 3*40 = 120 miles
-------
Equation:
distance = distance
8r = 120
rate = 15 mph (rate of the 1st train)
-----------------------------------------
Cheers,
Stan H.
---------------------